OpenStudy (anonymous):

Please, how do I go about this problem: d^2/dxdy (2x + 3y). How do I go about finding the derivative? Do I take the derivative of the function w.r.t to x then find the derivative of the answer w.r.t y?

OpenStudy (anonymous):

First take derivative with respect to x, then with respect to y. d^2/dxdy = d/dx (d/dy) = d/dy (d/dx) (but with curly d instead of regular d. remember, they are not the same thing!)

OpenStudy (waynex):

By the way, d^2/dxdy (2x + 3y) = d^2/dydx (2x + 3y). The order does not matter.

OpenStudy (anonymous):

It's worth noting though, that there exist some pathological examples when that is not the case. It happens when derivatives are not continous.